Are Japanese managers more motivated than American managers? A randomly selected group of each were administered the Sarnoff Survey of Attitudes Toward Life (SSATL), which measures motivation for upward mobility. The SSATL scores are summarized below.

American Japanese
Sample Size 211 100
Sample Mean SSATL Score 65.75 79.83
Sample Std. Dev. 11.07 6.41


What is the value of the test statistic?

-11.8092

-1.9677

-14.08

96.4471
Answer #1

standard error,
SE = sqrt(s1^2/n1 + s2^2/n2)
= sqrt(11.07^2/211 + 6.41^2/100)
= 0.9958

Test statistic,
t = (x1bar - x2bar)/SE
t = (65.75 - 79.83)/0.9958
t = -14.08

Option C